### v1

# Generic flatness of the cohomology of thickenings

Abstract.
Given a set of $m$ distinct points in projective space over a field, and $t$ a positive integer, a classical question asks for the least degree of a hypersurface that passes through each point with multiplicity at least $t$. Related to this, it remains unresolved whether there exists a dense open set of $m$-tuples of points for which this least degree is constant for each $t\geq 1$. We formulate a conjecture regarding the generic freeness of certain local cohomology modules that would answer this in the affirmative; we then prove this conjecture for up to $n+2$ points in projective $n$-space over a field. In a related direction, we prove a generic flatness result for the cohomology of thickenings of a closed subscheme of projective space over an integral domain.
2020 Mathematics Subject Classification:
Primary 13D45; Secondary 13A02, 13A30, 14F17.
Y.C.-R. was supported by NSF grant DMS 2502321, and A.K.S. by NSF grants DMS 2101671 and DMS 2349623.
1. Introduction
Let $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ be a projective scheme over a field $\mathbb{k}$. For each integer $t\geq 1$, the $t$-th thickening of $X$ is the closed subscheme $X_{t}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V\left(\mathscr{I}_{X}^{t}\right)$ in $\mathbb{P}_{\mathbb{k}}^{n}$, where $\mathscr{I}_{X}\subset\mathcal{O}_{\mathbb{P}_{\mathbb{k}}^{n}}$ is the ideal sheaf of $X$. The study of the asymptotic behavior of the cohomology of these thickenings has a long history in algebraic geometry and commutative algebra, with connections to deformation theory, Rees algebras, and local cohomology. More precisely, given a coherent sheaf $\mathcal{F}$ on $\mathbb{P}_{\mathbb{k}}^{n}$ that is flat along $X$, one is interested in asymptotically studying the induced projective system
[…]
for each cohomological index $i\geq 0$. We summarize some work in this direction:
(a) A foundational result of Hartshorne [Har66] shows that if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is smooth, where $\mathbb{k}$ has characteristic zero, then the cohomology of the symmetric powers of the conormal bundle exhibits strong vanishing properties; as a consequence, for each $i<\dim X$, the natural map
(1.0.1) $\begin{CD}\operatorname{H}^{i}\big(X_{t+1},\,\mathcal{O}_{X_{t+1}}\otimes_{\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{k}}}}\mathcal{F}\big)@>{}>{}>\operatorname{H}^{i}\big(X_{t},\,\mathcal{O}_{X_{t}}\otimes_{\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{k}}}}\mathcal{F}\big)\end{CD}$
is an isomorphism for all $t\gg 0$; see [Har66, Theorem 8.1 (b)] and [BBL+19, Theorem 2.8].
(b) This is extended in [BBL+19, Theorem 1.1] as follows: if $X\subset\mathbb{P}_{\mathbb{k}}^{n}$ is locally a complete intersection, where $\mathbb{k}$ is a field of characteristic zero, then, for each $i<\operatorname{codim}(\operatorname{Sing}X)$, the map () 1.0.1 is an isomorphism for all $t\gg 0$. In particular, if $X$ has at most isolated singular points, then () 1.0.1 is an isomorphism for all $i<\dim X$ and $t\gg 0$.
(c) A version of the Kodaira vanishing theorem for thickenings of locally complete intersection subvarieties of $\mathbb{P}_{\mathbb{k}}^{n}$, for $\mathbb{k}$ of characteristic zero, may be found in [BBL+19, Theorem 1.4]; an asymptotic version of the same holds in positive prime characteristic, [BBL+21, Theorem 1.2].
The present paper investigates a relative version of these phenomena over a Noetherian base ring; specifically, we study the generic flatness of the cohomology of thickenings of a closed subscheme in the following context: Let $A$ be a Noetherian domain, and $X\subset\mathbb{P}^{n}_{A}$ a closed subscheme. Consider the $t$-th thickening $X_{t}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V(\mathscr{I}_{X}^{t})$ of $X$ in $\mathbb{P}_{A}^{n}$, where $\mathscr{I}_{X}\subset\mathcal{O}_{\mathbb{P}_{A}^{n}}$ is the ideal sheaf of $X$. For $\mathcal{F}$ a coherent sheaf on $\mathbb{P}_{A}^{n}$ that is flat along $X$, a natural question then arises regarding the behavior of the cohomology groups
$\operatorname{H}^{i}\big(X_{t},\,\mathcal{O}_{X_{t}}\otimes_{\mathcal{O}_{\mathbb{P}_{A}^{n}}}\mathcal{F}\big)$
as $t$ varies: generic flatness guarantees that for a fixed $t$, the cohomology is flat over a dense open subset of $\operatorname{Spec}A$; however, it is not clear whether there is a single dense open subset that suffices for all $t\geq 1$. Our main result provides an affirmative answer to this when $X$ is smooth over $\operatorname{Spec}A$:
Theorem A.
Let $A$ be a Noetherian domain containing a field of characteristic zero. Let $X\subset\mathbb{P}^{n}_{A}$ be a closed subscheme that is smooth over $\operatorname{Spec}A$, and $X_{t}$ its $t$-th thickening, as defined above. Fix an integer $j$, and set $\mathcal{F}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{O}_{\mathbb{P}^{n}_{A}}(j)$. Then there exists a nonzero element $a\in A$ such that
$\operatorname{H}^{i}\big(X_{t},\,\mathcal{O}_{X_{t}}\otimes_{\mathcal{O}_{\mathbb{P}^{n}_{A}}}\mathcal{F}\big)\otimes_{A}A_{a}$
is a flat $A_{a}$-module for each integer $i\geq 0$, for all thickenings $t\geq 1$.
The proof of this relies on understanding the asymptotic behavior of the cohomology of powers of the ideal sheaf. Towards this, we use the notion of relatively ample vector bundles; while the theory of ample vector bundles on schemes over a field is well-established following Hartshorne [Har66], we require a relative version over an arbitrary Noetherian base scheme $S$. In Section 2, we develop the necessary theory of $f$-ample vector bundles and establish a fiberwise criterion for ampleness, where $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ is a projective morphism. Our primary example of an $f$-ample vector bundle is the normal bundle $\mathscr{N}_{X}$ associated with a closed immersion $X\lhook\joinrel\longrightarrow\mathbb{P}^{n}_{S}$, assuming that $X$ is smooth over the base $S$.
A key technical tool in our approach is a uniform vanishing result, Theorem 2.4, for the cohomology of the symmetric powers of the conormal bundle, that extends classical results of Hartshorne. In Section 3, we combine this vanishing result with a generic freeness result for local cohomology modules established in [CRS24]. More precisely, we exploit the following:
(a) In [CRS24, Theorem B], it was shown that the local cohomology modules $\operatorname{H}_{I}^{i}(R)$ are generically free over the base ring $A$, where $I$ is an ideal of $R\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}A[x_{0},\dots,x_{n}]$.
(b) Using a graded local duality argument, we obtain a stabilization result for the graded components of the modules $\operatorname{Ext}_{R}^{i}(R/I^{t},\,R)$, associated with the thickenings of $V(I)$ in $\mathbb{P}_{A}^{n}$.
(c) We deduce the flatness of the cohomology of thickenings by exploiting the natural identification
[…]
and the aforementioned stabilization result for Ext modules; see Section 3 and Section 3.
Our study of the generic flatness of cohomology modules is motivated by a classical question:
Question 1.1.
Given a set $X$ of $m$ distinct points in $\mathbb{P}_{\mathbb{k}}^{n}$, what is the minimum degree $\alpha_{t}(X)$ of a hypersurface in $\mathbb{P}_{\mathbb{k}}^{n}$ that passes through each point with multiplicity at least $t$?
The question gained significance following Nagata’s work on Hilbert’s 14-th problem [Nag59]; towards constructing his counterexample, Nagata proved that if $m\geq 16$ is a perfect square, then
$\alpha_{t}(X)\ >\ t\sqrt{m}.$
Substantial effort has since been put forth towards Section 1, though it remains elusive and difficult; cf. [Chu81, Dem82, EV83, BH10, Dum12, HH13, GHM13, Dum15, DTG17, FMX18, MSS18, BGHN22]. By a celebrated theorem of Zariski and Nagata, see [Zar49, Nag62], [Eis95, Theorem 3.14], when $\mathbb{k}$ is perfect, one has
$\alpha_{t}(X)\ =\ \alpha\big(I_{X}^{(t)}\big)\quad\text{ for all }t\geq 1,$
where $I_{X}$ is the defining ideal of $X$, the ideal $I_{X}^{(t)}$ is its $t$-th symbolic power, and for $I$ a homogeneous ideal, $\alpha(I)$ denotes the least degree of a homogeneous element of $I$.
Given an $m$-tuple $\mathbf{p}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\mathbf{p}_{1},\dots,\mathbf{p}_{m})$ of distinct points from $\mathbb{P}_{\mathbb{k}}^{n}$, set $X_{\mathbf{p}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{\mathbf{p}_{1},\dots,\mathbf{p}_{m}\}$ in $\mathbb{P}_{\mathbb{k}}^{n}$. To the best of our knowledge the following is completely open:
Question 1.2.
Does there exist a dense open subset $\mathcal{U}$ in $\left(\mathbb{P}_{\mathbb{k}}^{n}\right)^{m}$, such that the function
$\mathbf{p}\in\mathcal{U}\ \longmapsto\ \alpha_{t}\big(X_{\mathbf{p}}\big)$
is constant for all $t\geq 1$?
In other words, we ask whether the invariants $\alpha_{t}(X)$ are all constant for a general set of $m$ points in $\mathbb{P}_{\mathbb{k}}^{n}$. Again, a generic flatness argument yields such a dense open subset for a fixed $t$, but the challenge lies in finding a single dense open subset that works for all $t\geq 1$. We propose a conjecture that would settle the above question in the affirmative:
Conjecture 1.3.
Let $\mathbf{a}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left\{a_{ij}\mid 1\leq i\leq m,\,0\leq j\leq n\right\}$ be indeterminates over a field $\mathbb{k}$. Set $A\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbb{k}[\mathbf{a}]$, and let $R$ be the polynomial ring $A[x_{0},\dots,x_{n}]$. Consider the ideal
$I\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathfrak{A}_{1}\cap\dots\cap\mathfrak{A}_{m}$
of $R$, where $\mathfrak{A}_{i}$ is the ideal of size $2$ minors
$\mathfrak{A}_{i}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}I_{2}\begin{pmatrix}x_{0}&x_{1}&\cdots&x_{n}\\ a_{i0}&a_{i1}&\cdots&a_{in}\end{pmatrix}\quad\text{ for }1\leq i\leq m.$
In other words, $I$ is the defining ideal of a generic set of $m$ points in projective space. Set $\mathscr{R}(I)$ to be the Rees algebra $\bigoplus_{t\geq 0}I^{t}T^{t}\subset R[T]$. Then there exists a nonzero element $a\in A$ such that
$\operatorname{H}^{i}_{(x_{0},\dots,x_{n})}\left(\mathscr{R}(I)\right)\otimes_{A}A_{a}$
is a free $A_{a}$-module for each $i\geq 0$.
In Section 4 we prove a special case of this:
Theorem B.
Section 1 holds when $m\leq n+2$; in particular, the answer to Section 1 is affirmative for up to $n+2$ points in $\mathbb{P}_{\mathbb{k}}^{n}$.
2. Ample vector bundles relative to a base scheme
The notion of ample vector bundles on a scheme of finite type over a field was introduced by Hartshorne [Har66]; his main motivation was precisely to study the cohomology of thickenings as in () 1.0.1. For our purposes, we require a relative version of this notion over an arbitrary Noetherian base scheme. While this extension is possibly well-known to experts, we were unable to locate a reference that treats it in a form suitable for our needs. Accordingly, in this short section, we develop the necessary relative theory.
Let $S$ be a Noetherian scheme and $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ be a projective morphism. Let $\mathscr{E}$ be a vector bundle on $X$, i.e., a locally free sheaf of finite rank. Consider the projectivization
[…]
where we follow Grothendieck’s convention regarding projective bundles. Let
$g=f\circ\pi\colon\mathbf{P}(\mathscr{E})\ \longrightarrow\ S$
denote the composition of $f$ and $\pi$. By extending the notion of a relatively ample line bundle, see, e.g., [Sta25, Tag 01VG], we obtain the following definition:
Definition 2.1.
A vector bundle $\mathscr{E}$ on $X$ is $f$-ample if the canonical line bundle $\mathcal{O}_{\mathbf{P}(\mathscr{E})}(1)$ is $g$-ample.
Let $\kappa(s)$ denote the residue field at a point $s\in S$. Consider the fiber
$X_{s}\ \mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\ X\times_{S}\operatorname{Spec}(\kappa(s))\ \subset\ \mathbb{P}_{\kappa(s)}^{n}$
and the restriction vector bundle $\mathscr{E}_{s}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathscr{E}\otimes_{\mathcal{O}_{S}}\kappa(s)$ on $X_{s}$. Using the fiberwise criterion for the ampleness of line bundles, [Gro61b, Théorème 4.7.1] and [Gro66, Corollaire 9.6.4], we derive a corresponding fiberwise criterion for the ampleness of vector bundles:
Theorem 2.2.
The vector bundle $\mathscr{E}$ on $X$ is $f$-ample if and only if the restriction vector bundle $\mathscr{E}_{s}$ on $X_{s}$ is ample for each $s\in S$.
Proof.
The aforementioned fiberwise criterion implies that $\mathcal{O}_{\mathbf{P}(\mathscr{E})}(1)$ is $g$-ample if and only if
$\mathcal{O}_{\mathbf{P}(\mathscr{E}_{s})}(1)\ \cong\ \mathcal{O}_{\mathbf{P}(\mathscr{E})}(1)\otimes_{\mathcal{O}_{S}}\kappa(s)$
is ample for each $s\in S$. ∎
Next, assume that $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ is a smooth morphism. Our object of interest is the normal bundle to $X$ in $\mathbb{P}_{S}^{n}$, namely
[…]
Lemma 2.3.
If $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ is a smooth morphism, then the normal bundle $\mathscr{N}_{X}$ is $f$-ample.
Proof.
For each $s\in S$, [Har70, Proposition III.2.1] implies that the normal bundle to $X_{s}$ in $\mathbb{P}_{\kappa(s)}^{n}$, i.e., $\mathscr{N}_{X}\otimes_{\mathcal{O}_{S}}\kappa(s)$, is an ample vector bundle on $X_{s}$. Therefore $\mathscr{N}_{X}$ is $f$-ample by Theorem 2.2. ∎
Our main interest here lies in the following extension of a well-known vanishing result of Hartshorne, [Har66, Theorem 8.1]; see also [BBL+19, Theorem 2.8]. A key point in the vanishing result below is the existence of a stabilization index that is uniform across all fibers.
Theorem 2.4.
Let $S$ be a Noetherian scheme over the field of rational numbers, and $f\colon X\subset\mathbb{P}_{S}^{n}\longrightarrow S$ be a smooth morphism of relative dimension $d\geq 0$. Let $\mathcal{F}$ be a coherent sheaf on $X$ that is flat over $S$, and set $\mathcal{F}_{s}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{F}\otimes_{\mathcal{O}_{S}}\kappa(s)$ for $s\in S$. Then there exists an integer $t_{0}>0$ such that
$\operatorname{H}^{i}\left(X_{s},\,\operatorname{Sym}^{t}\big(\mathscr{I}_{X_{s}}/\mathscr{I}_{X_{s}}^{2}\big)\otimes_{\mathcal{O}_{X_{s}}}\mathcal{F}_{s}\right)\ =\ 0$
for all $s\in S$, $t\geq t_{0}$, and $i\neq d$.
Proof.
Let $s\in S$. By Serre duality, e.g., [Har77, §III.7], the cohomology group displayed above is dual to the cohomology group
$\operatorname{H}^{d-i}\Big(X_{s},\,\left(\operatorname{Sym}^{t}(\mathscr{I}_{X_{s}}/\mathscr{I}_{X_{s}}^{2})\right)^{\vee}\otimes_{\mathcal{O}_{X_{s}}}\mathcal{F}_{s}^{\vee}\otimes_{X_{s}}\omega_{X_{s}}\Big),$
[…]
see, e.g., [Eis95, Proposition A2.7]. Consider the coherent sheaf
$\mathcal{G}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{F}^{\vee}\otimes_{\mathcal{O}_{X}}\omega_{X/S}$
on $X$. After the above considerations, it suffices to find an integer $t_{0}>0$ such that
$\operatorname{H}^{j}\big(X_{s},\,\operatorname{Sym}^{t}(\mathscr{N}_{X})\otimes_{\mathcal{O}_{X}}\mathcal{G}\otimes_{\mathcal{O}_{S}}\kappa(s)\big)\ =\ 0$
for all $s\in S$, $t\geq t_{0}$, and $j\neq 0$. We now proceed to prove this claim.
Let $\mathscr{N}=\mathscr{N}_{X}$, $\pi\colon\mathbf{P}(\mathscr{N})\longrightarrow X$, and $g=f\circ\pi\colon\mathbf{P}(\mathscr{N})\longrightarrow S$. By Section 2, $\mathscr{N}$ is an $f$-ample vector bundle on $X$, so $\mathcal{O}_{\mathbf{P}(\mathscr{N})}(1)$ is a $g$-ample line bundle on $\mathbf{P}(\mathscr{N})$. Consider the Grothendieck spectral sequence
$E_{2}^{p,q}\ =\ R^{p}f_{*}\left(R^{q}\pi_{*}\big(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\big)\right)\ \Longrightarrow\ R^{p+q}g_{*}\left(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\right).$
We have the isomorphism $\pi_{*}\big(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\big)\ \cong\ \operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}$ and the vanishing
$R^{q}\pi_{*}\left(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\right)\ =\ 0\quad\text{ for all }q\geq 1;$
see [Har66, Lemma 3.1], [Har77, Exercise III.8.4]. Hence, the above spectral sequence degenerates and yields the isomorphism
$R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)\ \cong\ R^{j}g_{*}\left(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\right)\quad\text{ for all }j\geq 0.$
Since $\mathcal{O}_{\mathbf{P}(\mathscr{N})}(1)$ is $g$-ample, there is an integer $t_{0}>0$ such that
$R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)\ \cong\ R^{j}g_{*}\left(\mathcal{O}_{\mathbf{P}(\mathscr{N})}(t)\otimes_{\mathcal{O}_{\mathbf{P}(\mathscr{N})}}\pi^{*}(\mathcal{G})\right)\ =\ 0$
for all $j\geq 1$ and $t\geq t_{0}$, see [Sta25, Tag 02O1], [Har77, Theorem III.8.8].
Fix $t\geq t_{0}$. Consider the natural base change map
$\begin{CD}\varphi^{j}(s)\colon R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)\otimes_{\mathcal{O}_{S}}\kappa(s)@>{}>{}>\operatorname{H}^{j}\left(X_{s},\,\operatorname{Sym}^{t}(\mathscr{N}_{s})\otimes_{\mathcal{O}_{X_{s}}}\mathcal{G}_{s}\right).\end{CD}$
The cohomology and base change theorem, [Har77, Theorem III.12.11], [Gro63, §7.7], [Mum08, §5], tells us that: (a) $\varphi^{j}(s)$ is an isomorphism if it is surjective; and (b) if $\varphi^{j}(s)$ is surjective, then $\varphi^{j-1}(s)$ is surjective if and only if $R^{j}f_{*}\left(\operatorname{Sym}^{t}(\mathscr{N})\otimes_{\mathcal{O}_{X}}\mathcal{G}\right)$ is locally free in a neighborhood of $s$.
Finally, by descending induction on $j\leq d$, since $\varphi^{d}(s)$ is surjective, we obtain the vanishing
$\operatorname{H}^{j}\left(X_{s},\,\operatorname{Sym}^{t}(\mathscr{N}_{s})\otimes_{\mathcal{O}_{X_{s}}}\mathcal{G}_{s}\right)\ =\ 0$
for all $j\geq 1$, see [Mum08, §5, Corollary 4]. This completes the proof of the theorem. ∎
3. Generic flatness in the smooth setting
In this section, we complete the proof of Theorem A, i.e., under a smoothness hypothesis, we prove that there exists a flatness locus for the cohomology of all thickenings. The main tools are the vanishing result, Theorem 2.4, and the generic freeness result for local cohomology modules stated in Theorem 3.3.
We fix the notation that we will use throughout this section. Let $A$ be a Noetherian domain containing a field of characteristic zero. Let $R\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}A[x_{0},\dots,x_{n}]$ be a standard graded polynomial ring over $A$, i.e., $\deg(x_{i})=1$, and $\deg(a)=0$ for $a\in A$. Set $\mathfrak{M}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(x_{0},\dots,x_{n})$ and $\mathbb{P}_{A}^{n}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\operatorname{Proj}(R)$.
Let $Q\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\operatorname{Frac}A$ be the field of fractions of $A$. For $\mathfrak{p}\in\operatorname{Spec}A$, let $\kappa(\mathfrak{p})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ be the residue field at $\mathfrak{p}$, and set
$R(\mathfrak{p})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R\otimes_{A}\kappa(\mathfrak{p})\ \cong\ \kappa(\mathfrak{p})[x_{0},\dots,x_{n}].$
For $M$ a finitely generated graded $R$-module, set $M(\mathfrak{p})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}M\otimes_{A}\kappa(\mathfrak{p})$.
Let $X\subset\mathbb{P}_{A}^{n}$ be a closed subscheme with ideal sheaf $\mathscr{I}_{X}\subset\mathcal{O}_{\mathbb{P}_{A}^{n}}$. For $\mathfrak{p}\in\operatorname{Spec}A$, consider the fiber
$X_{\mathfrak{p}}\ \mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\ X\times_{\operatorname{Spec}A}\operatorname{Spec}(\kappa(\mathfrak{p}))\ \subset\ \mathbb{P}_{\kappa(\mathfrak{p})}^{n}.$
For $t\geq 1$, the $t$-th thickening of $X$ in $\mathbb{P}_{A}^{n}$ is the closed subscheme of $\mathbb{P}_{A}^{n}$ determined by the $t$-th power of the ideal sheaf $\mathscr{I}_{X}$ of $X$, namely
$X_{t}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V\left(\mathscr{I}_{X}^{t}\right)\ \subset\ \mathbb{P}_{A}^{n}.$
[…]
The following proposition yields a version of the graded local duality theorem relative to our Noetherian base ring $A$. Results of this type have also appeared in [HR76, Smi18, CCRS22, CR23].
Proposition 3.1.
Let $R=A[x_{0},\dots,x_{n}]$ be a standard graded polynomial ring over $A$, and $M$ a finitely generated graded $R$-module that is flat over $A$. Fix an integer $j$. Then the following are equivalent:
  1.  (a)
    ${\operatorname{H}_{\mathfrak{M}}^{i}(M)}_{j}$ is a flat $A$-module for each $i\geq 0$.
  2.  (b)
    ${\operatorname{Ext}_{R}^{i}(M,\,R)}_{-j-n-1}$ is a flat $A$-module for each $i\geq 0$.
  3.  (c)
    The function $\operatorname{Spec}A\longrightarrow\mathbb{Z}$ where $\mathfrak{p}\longmapsto\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{H}_{\mathfrak{M}}^{i}\left(M(\mathfrak{p})\right)}_{j}$ is constant for each $i\geq 0$.
  4.  (d)
    The function $\operatorname{Spec}A\longrightarrow\mathbb{Z}$ where $\mathfrak{p}\longmapsto\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(M(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{-j-n-1}$ is constant for each $i\geq 0$.
Moreover, when these equivalent conditions are satisfied, we have the base change isomorphisms
$\begin{CD}{\operatorname{H}_{\mathfrak{M}}^{i}\left(M\right)}_{j}\otimes_{A}B@>{\cong}>{}>{\operatorname{H}_{\mathfrak{M}}^{i}\left(M\otimes_{A}B\right)}_{j}\quad\text{ and }\quad{\operatorname{Ext}_{R}^{i}\left(M,\,R\right)}_{j}\otimes_{A}B@>{\cong}>{}>{\operatorname{Ext}_{R_{B}}^{i}\left(M\otimes_{A}B,\,R_{B}\right)}_{j}\end{CD}$
for each $i\geq 0$, where $B$ is an arbitrary $A$-algebra, and $R_{B}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R\otimes_{A}B$.
Proof.
The equivalence (a) $\iff$ (b) follows from [CR23, Theorem A]; one can restrict this result to the $j$-th graded component, see also the proof of [CR23, Proposition 2.11]. The base change isomorphisms may also be obtained from [CR23, Theorem A].
The equivalence (a) $\iff$ (c) follows from classical results on cohomological functors, see for example [Gro63, Théorème 7.6.9] or [Har77, Corollary 12.9]. The equivalence (c) $\iff$ (d) may be obtained by applying the graded local duality theorem [BH98, Theorem 3.6.19] to the module $M(\mathfrak{p})$. ∎
For our purposes, one of the principal applications of Section 3 is the following corollary; this enables us to apply the vanishing proven in Theorem 2.4.
Corollary 3.2.
[…]
  1.  (a)
    $M$ is $A$-flat,
  2.  (b)
    $M_{0}=0$,
  3.  (c)
    $\operatorname{H}^{i}\big(X_{\mathfrak{p}},\,\mathcal{F}_{\mathfrak{p}}\big)=0$ for all $i\neq d$ and $\mathfrak{p}\in\operatorname{Spec}A$.
Then ${\operatorname{Ext}_{R}^{n-d}(M,\,R)}_{-n-1}$ is $A$-flat, and ${\operatorname{Ext}_{R}^{n-i}(M,\,R)}_{-n-1}=0$ for all $i\neq d$. Moreover, we have
${\operatorname{Ext}_{R}^{i}(M,\,R)}_{-n-1}\otimes_{A}B\ \cong\ {\operatorname{Ext}_{R_{B}}^{i}(M_{B},\,R_{B})}_{-n-1}$
for each $i\geq 0$, where $B$ is an arbitrary $A$-algebra, $R_{B}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R\otimes_{A}B$, and $M_{B}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}M\otimes_{A}B$.
Proof.
We have an exact sequence
[…]
and an isomorphism
$\operatorname{H}_{\mathfrak{M}}^{i+1}\left(M(\mathfrak{p})\right)\,\cong\,\bigoplus_{j\in\mathbb{Z}}\operatorname{H}^{i}\big(X_{\mathfrak{p}},\,\mathcal{F}_{\mathfrak{p}}(j)\big)\quad\text{ for all }i\geq 1.$
From the assumptions, we obtain
${\operatorname{H}_{\mathfrak{M}}^{i+1}\left(M\left(\mathfrak{p}\right)\right)}_{0}\ =\ 0\quad\text{ for all }i\neq d.$
Moreover, since the function
$\mathfrak{p}\in\operatorname{Spec}A\ \longmapsto\ \chi\left(\mathcal{F}_{\mathfrak{p}}\right)\ =\ \sum_{i\geq 0}(-1)^{i}\operatorname{rank}_{\kappa(\mathfrak{p})}\operatorname{H}^{i}(X_{\mathfrak{p}},\,\mathcal{F}_{\mathfrak{p}})$
is constant, [Mum08, §5], we conclude that $\mathfrak{p}\in\operatorname{Spec}A\longmapsto\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{H}_{\mathfrak{M}}^{d+1}\left(M\left(\mathfrak{p}\right)\right)}_{0}$ is constant as well. The result now follows from Section 3. ∎
We next recall a result on the generic freeness of local cohomology modules from [CRS24]:
Theorem 3.3 ([CRS24, Theorem B]).
Let $A$ be a Noetherian domain containing a field of characteristic zero, and $I$ an ideal of $R=A[x_{0},\dots,x_{n}]$. Then there is a nonzero element $a\in A$ such that
$\operatorname{H}_{I}^{i}(R)\otimes_{A}A_{a}$
is $A_{a}$-free for each $i\geq 0$.
Remark 3.4.
When $\operatorname{H}_{I}^{i}(R)\otimes_{A}A_{a}$ is $A_{a}$-flat for each $i\geq 0$, one has base change isomorphisms
$\begin{CD}\operatorname{H}_{I}^{i}(R)\otimes_{A}B@>{\cong}>{}>\operatorname{H}_{I}^{i}(R\otimes_{A}B)\end{CD}$
for each $i\geq 0$, where $B$ is an arbitrary $A_{a}$-algebra.
The next proposition provides a crucial stabilization argument; it allows us to exploit the generic flatness of local cohomology, Theorem 3.3, to study Ext modules:
Proposition 3.5.
Let $I\subset R$ be a homogeneous ideal and $X=V(I)\subset\mathbb{P}_{A}^{n}$ be the closed subscheme defined by $I$. Fix an integer $j\in\mathbb{Z}$. Assume that:
  1.  (a)
    The natural projection $X\subset\mathbb{P}_{A}^{n}\longrightarrow\operatorname{Spec}A$ is a smooth morphism of relative dimension $d\geq 0$,
  2.  (b)
    $I^{t}/I^{t+1}$ is $A$-flat for each $t\geq 0$,
  3.  (c)
    $\operatorname{H}_{I}^{i}(R)$ is $A$-flat for each $i\geq 0$.
Then there is an integer $t_{0}>0$, such that the natural map
$\begin{CD}\psi_{B,t}^{i}\colon{\operatorname{Ext}_{R_{B}}^{i}\left(R_{B}/I^{t}R_{B},\,R_{B}\right)}_{-j-n-1}@>{}>{}>{\operatorname{Ext}_{R_{B}}^{i}\left(R_{B}/I^{t+1}R_{B},\,R_{B}\right)}_{-j-n-1}\end{CD}$
is injective for each $t\geq t_{0}$ and $i\geq 0$, where $B$ is an arbitrary $A$-algebra, and $R_{B}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R\otimes_{A}B$.
Proof.
Set $k\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}-j-n-1$, and let $B$ be an $A$-algebra. Taking $\mathcal{F}=\mathcal{O}_{X}(j)$ in Theorem 2.4, there exists an integer $t_{0}>0$ such that
$\operatorname{H}^{i}\Big(X_{\mathfrak{p}},\,\operatorname{Sym}^{t}\big(\mathscr{I}_{X_{\mathfrak{p}}}/\mathscr{I}_{X_{\mathfrak{p}}}^{2}\big)(j)\Big)\ =\ 0$
for all $\mathfrak{p}\in\operatorname{Spec}A$, $t\geq t_{0}$, and $i\neq d$. After possibly increasing $t_{0}$, we may assume that $\left[I^{t}/I^{t+1}\right]_{j}=0$ for all $t\geq t_{0}$. Applying Section 3 to $M=\left(I^{t}/I^{t+1}\right)(j)$ and $\widetilde{M}\cong\operatorname{Sym}^{t}(\mathscr{I}_{X}/\mathscr{I}_{X}^{2})(j)$ yields
${\operatorname{Ext}_{R_{B}}^{n-i}\left(I^{t}R_{B}/I^{t+1}R_{B},\,R_{B}\right)}_{k}\ =\ 0$
for all $t\geq t_{0}$ and $i\neq d$.
For $t\geq t_{0}$, consider the exact sequence
$\begin{CD}0@>{}>{}>I^{t}R_{B}/I^{t+1}R_{B}@>{}>{}>R_{B}/I^{t+1}R_{B}@>{}>{}>R_{B}/I^{t}R_{B}@>{}>{}>0.\end{CD}$
Setting $c\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}n-d$, the associated exact sequence of Ext modules gives:
[…]
It only remains to deal with the case $i=c+1$. For each $t\geq t_{0}$, we have an exact sequence
$\begin{CD}0@>{}>{}>K_{t}@>{}>{}>{\operatorname{Ext}_{R}^{c+1}(R/I^{t_{0}},\,R)}_{k}@>{}>{}>{\operatorname{Ext}_{R}^{c+1}(R/I^{t},\,R)}_{k}@>{}>{}>0,\end{CD}$
with $K_{t}$ denoting the kernel. By the Noetherian hypothesis, the ascending sequence
$K_{t_{0}}\;\subseteq\;K_{t_{0}+1}\;\subseteq\;K_{t_{0}+2}\;\subseteq\;\dots$
inside ${\operatorname{Ext}_{R}^{c+1}(R/I^{t_{0}},\,R)}_{k}$ must be eventually constant. Thus, after possibly increasing $t_{0}$, we may assume that $\psi_{A,t}^{c+1}$ is an isomorphism for all $t\geq t_{0}$. Hence we may assume that
${\operatorname{Ext}_{R}^{i}(R/I^{t},\,R)}_{k}\ \cong\ {\operatorname{H}_{I}^{i}(R)}_{k}\quad\text{ for all $i\geq c+1$ and $t\geq t_{0}$;}$
in particular, by assumption, these modules are locally free over $A$.
Fix $\mathfrak{p}\in\operatorname{Spec}A$ and $t\geq t_{0}$, and write $R_{\mathfrak{p}}=R\otimes_{A}A_{\mathfrak{p}}$ and $B_{\mathfrak{p}}=B\otimes_{A}A_{\mathfrak{p}}$. Consider the natural base change map
$\begin{CD}b^{i}(\mathfrak{p})\colon{\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}\kappa(\mathfrak{p})@>{}>{}>{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(R(\mathfrak{p})/I^{t}R(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{k}.\end{CD}$
The property of exchange for local Ext’s, [AK80, Theorem 1.9], [LK79, Theorem A.5], implies that, if $b^{i}(\mathfrak{p})$ is surjective, then:
(a) $b^{i}(\mathfrak{p})$ is an isomorphism;
(b) the base change map below is also an isomorphism
$\begin{CD}{\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}B_{\mathfrak{p}}@>{}>{}>{\operatorname{Ext}_{R_{B}}^{i}\left(R_{B}/I^{t}R_{B},\,R_{B}\right)}_{k}\otimes_{A}A_{\mathfrak{p}};\end{CD}$
(c) $b^{i-1}(\mathfrak{p})$ is surjective if and only if ${\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}$ is $A_{\mathfrak{p}}$-free.
Since we know that ${\operatorname{Ext}_{R_{\mathfrak{p}}}^{i}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}$ is $A_{\mathfrak{p}}$-free for all $i\geq c+1$, by descending induction on $i$ (starting with $i=n+1$), we get the isomorphism
$\begin{CD}{\operatorname{Ext}_{R_{\mathfrak{p}}}^{c+1}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}B_{\mathfrak{p}}@>{\cong}>{}>{\operatorname{Ext}_{R_{B}}^{c+1}\left(R_{B}/I^{t}R_{B},\,R_{B}\right)}_{k}\otimes_{A}A_{\mathfrak{p}}.\end{CD}$
Therefore, we obtain the isomorphisms
${\operatorname{Ext}_{R_{B}}^{c+1}\left(R_{B}/I^{t}R_{B},\,R_{B}\right)}_{k}\otimes_{A}A_{\mathfrak{p}}$ $\ \cong\ {\operatorname{Ext}_{R_{\mathfrak{p}}}^{c+1}\left(R_{\mathfrak{p}}/I^{t}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}B_{\mathfrak{p}}$
$\ \cong\ {\operatorname{H}_{I}^{c+1}\left(R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}B_{\mathfrak{p}}$
$\ \cong\ {\operatorname{Ext}_{R_{\mathfrak{p}}}^{c+1}\left(R_{\mathfrak{p}}/I^{t+1}R_{\mathfrak{p}},\,R_{\mathfrak{p}}\right)}_{k}\otimes_{A_{\mathfrak{p}}}B_{\mathfrak{p}}$
$\ \cong\ {\operatorname{Ext}_{R_{B}}^{c+1}\left(R_{B}/I^{t+1}R_{B},\,R_{B}\right)}_{k}\otimes_{A}A_{\mathfrak{p}}$
for all $t\geq t_{0}$. Finally, as $\mathfrak{p}\in\operatorname{Spec}A$ was arbitrary, we get the isomorphism
$\begin{CD}{\operatorname{Ext}_{R_{B}}^{c+1}(R_{B}/I^{t}R_{B},\,R_{B})}_{k}@>{\cong}>{}>{\operatorname{Ext}_{R_{B}}^{c+1}(R_{B}/I^{t+1}R_{B},\,R_{B})}_{k}\end{CD}$
for all $t\geq t_{0}$. This concludes the prove of the proposition. ∎
The following lemma yields a flatness criterion, assuming Theorem 3.3 and Section 3.
Lemma 3.6 (Valuative criterion for the flatness of cohomology).
Let $R=A[x_{0},\dots,x_{n}]$ be a standard graded polynomial ring over a Noetherian domain $A$, and $I\subset R$ a homogeneous ideal. Fix $j\in\mathbb{Z}$ and $t\geq 1$. Assume that:
  1.  (a)
    For each $A$-algebra $V$ that is a discrete valuation ring, and for each $i\geq 0$, the natural map
    $\begin{CD}{\operatorname{Ext}_{R_{V}}^{i}\left(R_{V}/I^{t}R_{V},\,R_{V}\right)}_{-j-n-1}@>{}>{}>{\operatorname{H}_{I}^{i}(R_{V})}_{-j-n-1}\end{CD}$
    is injective, where $R_{V}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R\otimes_{A}V$,
  2.  (b)
    $R/I^{t}$ is $A$-flat,
  3.  (c)
    $\operatorname{H}_{I}^{i}(R)$ is $A$-flat for each $i\geq 0$.
Then ${\operatorname{H}_{\mathfrak{M}}^{i}(R/I^{t})}_{j}$ is $A$-flat for each $i\geq 0$.
Proof.
Set $k\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}-j-n-1$. By Section 3, it suffices to show that
$\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(R(\mathfrak{p})/I^{t}R(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{k}\ =\ \operatorname{rank}_{Q}{\operatorname{Ext}_{R}(R/I^{t},\,R)}_{k}\otimes_{A}Q$
for all $\mathfrak{p}\in\operatorname{Spec}A$, where $Q\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\operatorname{Frac}A$.
Fix $\mathfrak{p}\in\operatorname{Spec}A$. By [Har77, Exercise II.4.11] or [Gro61a, Proposition 7.1.7], there exists a discrete valuation ring $(V,\mathfrak{n})$ of $Q$ that dominates $A_{\mathfrak{p}}$, i.e., $A_{\mathfrak{p}}\subset V$ and $\mathfrak{p}A_{\mathfrak{p}}=\mathfrak{n}\cap A_{\mathfrak{p}}$. To simplify notation, set $\kappa(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}V/\mathfrak{n}$ and $R_{V}(\mathfrak{n})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}R_{V}\otimes_{V}\kappa(\mathfrak{n})\cong R_{V}/\mathfrak{n}R_{V}$. Then $\operatorname{H}_{I}^{i}(R_{V})\cong\operatorname{H}_{I}^{i}(R)\otimes_{A}V$ is flat over $V$, equivalently torsionfree over $V$, for each $i\geq 0$. Therefore the injective map
${\operatorname{Ext}_{R_{V}}^{i}\left(R_{V}/I^{t}R_{V},\,R_{V}\right)}_{k}\ \lhook\joinrel\longrightarrow\ {\operatorname{H}_{I}^{i}(R_{V})}_{k}$
from (a) implies that each ${\operatorname{Ext}_{R_{V}}^{i}\left(R_{V}/I^{t}R_{V},\,R_{V}\right)}_{k}$ is also torsionfree over $V$, hence flat over $V$. By applying Section 3 to $R_{V}/I^{t}R_{V}$ over $V$, we obtain
$\operatorname{rank}_{\kappa(\mathfrak{n})}{\operatorname{Ext}_{R_{V}(\mathfrak{n})}^{i}\left(R_{V}(\mathfrak{n})/I^{t}R_{V}(\mathfrak{n}),\,R_{V}(\mathfrak{n})\right)}_{k}$ $\ =\ \operatorname{rank}_{\kappa(\mathfrak{n})}{\operatorname{Ext}_{R_{V}}^{i}(R_{V}/I^{t}R_{V},\,R_{V})}_{k}\otimes_{V}\kappa(\mathfrak{n})$
$\ =\ \operatorname{rank}_{Q}{\operatorname{Ext}_{R}^{i}(R/I^{t},\,R)}_{k}\otimes_{A}Q.$
Considering the field extension $\kappa(\mathfrak{p})\lhook\joinrel\longrightarrow\kappa(\mathfrak{n})$, we have isomorphisms
$R_{V}(\mathfrak{n})\ \cong\ (R\otimes_{A}V)\otimes_{V}\kappa(\mathfrak{n})\ \cong\ R\otimes_{A}\kappa(\mathfrak{n})\ \cong\ (R\otimes_{A}\kappa(\mathfrak{p}))\otimes_{\kappa(\mathfrak{p})}\kappa(\mathfrak{n})\ \cong\ R(\mathfrak{p})\otimes_{\kappa(\mathfrak{p})}\kappa(\mathfrak{n}).$
Thus, we have the isomorphism
$\operatorname{Ext}_{R_{V}(\mathfrak{n})}^{i}\big(R_{V}(\mathfrak{n})/I^{t}R_{V}(\mathfrak{n}),\,R_{V}(\mathfrak{n})\big)\ \cong\ \operatorname{Ext}_{R(\mathfrak{p})}^{i}\big(R(\mathfrak{p})/I^{t}R(\mathfrak{p}),\,R(\mathfrak{p})\big)\otimes_{\kappa(\mathfrak{p})}\kappa(\mathfrak{n}).$
Combining everything, we obtain
$\operatorname{rank}_{\kappa(\mathfrak{p})}{\operatorname{Ext}_{R(\mathfrak{p})}^{i}\left(R(\mathfrak{p})/I^{t}R(\mathfrak{p}),\,R(\mathfrak{p})\right)}_{k}\ =\ \operatorname{rank}_{Q}{\operatorname{Ext}_{R}(R/I^{t},\,R)}_{k}\otimes_{A}Q,$
which completes the proof of the lemma. ∎
We are now ready to prove the main result:
Proof of Theorem A.
Let $I\subset R$ be a homogeneous ideal defining $X$ in $\mathbb{P}_{A}^{n}$. By generic freeness, e.g., [Mat89, §24], there exists a nonzero element $a^{\prime}$ in $A$ such that
$\operatorname{gr}_{I}(R)\otimes_{A}A_{a^{\prime}}\ =\ \bigoplus_{t\geq 0}\,I^{t}/I^{t+1}\otimes_{A}A_{a^{\prime}}$
is $A_{a^{\prime}}$-free. By Theorem 3.3, there exists a nonzero element $a^{\prime\prime}$ in $A$ such that $\operatorname{H}_{I}^{i}(R)\otimes_{A}A_{a^{\prime\prime}}$ is $A_{a^{\prime\prime}}$-free for each $i\geq 0$. To simplify notation, substitute $A$ by the localization $A_{a^{\prime}a^{\prime\prime}}$. Set k=jn1k=-j-n-1. Then Section 3 yields an integer $t_{0}>0$ such that the natural map
$\begin{CD}{\operatorname{Ext}_{R_{V}}^{i}\left(R_{V}/I^{t}R_{V},\,R_{V}\right)}_{k}@>{}>{}>{\operatorname{H}_{I}^{i}\left(R_{V}\right)}_{k}\end{CD}$
is injective for all $t\geq t_{0}$ and $i\geq 0$ and for any $A$-algebra $V$ that is a discrete valuation ring. From Section 3, we get that ${\operatorname{H}_{\mathfrak{M}}^{i}\left(R/I^{t}\right)}_{j}$ is $A$-flat for all $i\geq 0$ and $t\geq t_{0}$. After possibly increasing $t_{0}$, we may also assume that $\left[I^{t}/I^{t+1}\right]_{j}=0$ for all $t\geq t_{0}$. This implies that
$\operatorname{H}^{i}\big(X_{t},\,\mathcal{O}_{X_{t}}\otimes_{\mathcal{O}_{\mathbb{P}_{A}^{n}}}\mathcal{F}\big)$
is $A$-flat for all $i\geq 0$ and $t\geq t_{0}$, cf. proof of Section 3. Finally, by generic freeness, there exists a nonzero element $a\in A$ such that
$\operatorname{H}^{i}\big(X_{t},\,\mathcal{O}_{X_{t}}\otimes_{\mathcal{O}_{\mathbb{P}_{A}^{n}}}\mathcal{F}\big)\otimes_{A}A_{a}$
is $A_{a}$-free for all $1\leq t<t_{0}$ and $i\geq 0$. This concludes the proof of the theorem. ∎
4. Ideals of points
We return to the motivating conjecture for our work on the generic freeness of local cohomology:
Proof of Theorem B.
The ideals $\mathfrak{A}_{1},\dots\mathfrak{A}_{m}$, as defined in Section 1, are unchanged by column operations on the matrix
$\begin{pmatrix}x_{0}&x_{1}&\cdots&x_{n}\\ a_{10}&a_{11}&\cdots&a_{1n}\\ a_{20}&a_{21}&\cdots&a_{2n}\\ \vdots&\vdots&&\vdots\\ a_{m0}&a_{m1}&\cdots&a_{mn}\end{pmatrix}$.
Note that $\mathfrak{A}_{i}$ is the ideal of size $2$ minors of the submatrix given by the first and $(i+1)$-th rows. After inverting the element $a_{10}\in A$, we may subtract multiples of the first column from the other columns, so that the other entries in that row are zero. Next invert the element diagonally below $a_{10}$, and subtract suitable multiples of the second column from the others, so as to obtain zeros elsewhere in that row; proceed in this manner. Let $a$ denote the product of the elements that have been inverted. For notational simplicity, assume for the moment that $m=n+2$; the resulting matrix then has the form
$\begin{pmatrix}y_{0}&y_{1}&y_{2}&\cdots&y_{n}\\ a_{1}&0&0&\cdots&0\\ 0&a_{2}&0&\cdots&0\\ 0&0&a_{3}&&0\\ \vdots&\vdots&&\ddots\\ 0&0&0&&a_{n+1}\\ b_{0}&b_{1}&b_{2}&\cdots&b_{n}\end{pmatrix}$,
where $a_{i},b_{i}\in A_{a}$, and $y_{0},\dots,y_{n}$ are independent linear forms in the indeterminates $x_{0},\dots,x_{n}$. Specifically, the ideal $(x_{0},\dots,x_{n})R_{a}$ agrees with the ideal $(y_{0},\dots,y_{n})R_{a}$. After replacing the element $a$, we may also assume that the $b_{i}$ are units in $A_{a}$; dividing the columns by $b_{0},\dots,b_{n}$ respectively, and renaming the elements $a_{i}$ and $y_{i}$, the matrix takes the form
$\begin{pmatrix}y_{0}&y_{1}&y_{2}&\cdots&y_{n}\\ a_{1}&0&0&\cdots&0\\ 0&a_{2}&0&\cdots&0\\ 0&0&a_{3}&&0\\ \vdots&\vdots&&\ddots\\ 0&0&0&&a_{n+1}\\ 1&1&1&\cdots&1\end{pmatrix}$.
Note that
$\mathfrak{A}_{i}R_{a}\ =\ (y_{0},\ \dots,\ \widehat{y}_{i-1},\ \dots,\ y_{n})R_{a}\quad\text{ for }1\leq i\leq n+1,$
and that
$\mathfrak{A}_{n+2}R_{a}\ =\ (y_{1}-y_{0},\ y_{2}-y_{0},\ \dots,\ y_{n}-y_{0})R_{a}.$
In the case that $m<n+2$, the generators of the ideals $\mathfrak{A}_{1}R_{a},\dots,\mathfrak{A}_{m}R_{a}$ remain as displayed above, while the other ideals $\mathfrak{A}_{i}R_{a}$ do not occur. In either case, $\mathfrak{A}_{1}R_{a},\dots,\mathfrak{A}_{m}R_{a}$ are extended from the polynomial ring $P\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbb{k}[y_{0},\dots,y_{n}]$; set $J$ to be the intersection of the corresponding ideals in $P$, in which case
$IR_{a}\ =\ \mathfrak{A}_{1}R_{a}\cap\dots\cap\mathfrak{A}_{m}R_{a}\ =\ JR_{a}.$
Set $\mathscr{R}(J)$ to be the Rees algebra $\bigoplus_{t\geq 0}J^{t}T^{t}\subset P[T]$. For each $i\geq 0$, the local cohomology module
$\operatorname{H}^{i}_{(y_{0},\dots,y_{n})}\left(\mathscr{R}(J)\right)$
is, of course, free over the field $\mathbb{k}$, but then
$\operatorname{H}^{i}_{(y_{0},\dots,y_{n})}\left(\mathscr{R}(J)\right)\otimes_{\mathbb{k}}A_{a}\ \cong\ \operatorname{H}^{i}_{(x_{0},\dots,x_{n})}\left(\mathscr{R}(I)\right)\otimes_{A}A_{a}$
is a free $A_{a}$-module. ∎